# JigsawComm: Joint Semantic Feature Encoding and Transmission for Communication-Efficient Cooperative Perception

### v1

@@ 1 Introduction
Figure 1: A sample scenario from OPV2V dataset (left). Under the same bandwidth limit, prior work [hu2022where2comm] (mid) does not utilize bandwidth efficiently, leading to missed detection and reduced accuracy. JigsawComm (right) maximizes every bit’s contribution, enabling complete and accurate detection. Quantitative evaluation and discussion on the ‘cost’ of redundancy can be found in Sec. 4 and Sec. 5, respectively.
@@
Despite this promise, CP faces a formidable challenge in communication efficiency. The available V2X bandwidth is severely limited (up to $20$ MHz for both DSRC [FCC_DSRC] and C-V2X [CV2X_ETSI_EN_303_613]), while the volume of sensor data generated is prohibitively large (e.g. 4 MB per LiDAR frame [f_cooper]). Exchanging raw sensor data is therefore infeasible while satisfying the stringent real-time perception requirements of autonomous driving (e.g. <<100 ms [ADConstraints]). As previous research also demonstrated that exchanging local detection results alone is suboptimal [xu2022opv2v], the field has largely converged on the intermediate fusion framework—exchanging and fusing Bird’s-Eye-View (BEV) features—as the most effective paradigm [xu2022opv2v]. Even so, transmitting dense deep features from multiple agents still exhausts V2X bandwidth.
@@
The key challenge stems from the high degree of data redundancy present in these shared features, which can be classified into two forms: (i) Intra-agent redundancy, where features waste bits on semantically irrelevant information; and (ii) Inter-agent redundancy, where multiple agents transmit overlapping and unnecessary information. As shown in Fig. 1, under limited bandwidth, such inefficient communication not only leads to reduced detection coverage but can even degrade perception accuracy when lower quality data dilutes the utility of the better ones after fusion.
@@
Existing work on communication-efficient CP fails to solve this. They typically adopt one of two strategies: (i) Compression methods, such as encoder-decoders [xu2022opv2v] or codebooks [hu2024communication, zhao2025quantv2x]. Such methods either only reduce the data size but still transmit the entire feature map, failing to address redundancy, or overly restrict the representation power to offline-trained dictionaries, undermining generalizability; (ii) Message selection methods [liu2020when2com, hu2022where2comm, hu2024communication, xu2025cosdh], choosing ‘critical’ regions to transmit but relying on simple local heuristics (e.g. local detection confidence) that are poor proxies for the final cooperative task. Crucially, both approaches ignore cross-agent redundancy. As a result, they still require significant bandwidth and are not scalable: their communication cost grows linearly with the number of agents. The problem of designing a truly communication-efficient, scalable, and accurate CP system remains open.
@@
To design such a CP system, we argue that it must maximize the contribution of every transmitted bit towards the final perception goal. This requires that only essential, semantically meaningful, and non-redundant information is transmitted. Therefore, the feature encoding (local compression) and transmission (message selection) algorithms must be jointly designed to realize this potential. This joint design is non-trivial. Removing semantic, cross-agent redundancy is a challenging ‘chicken-and-egg’ problem: an agent cannot determine if its features are redundant without first knowing what other agents intend to transmit.
@@
To address these gaps, we propose JigsawComm, an end-to-end semantic-aware CP framework that learns to “assemble the puzzle” of multi-agent feature transmission. Our approach is based on the joint semantic feature encoding and transmission problem that we formulate. First, a regularized feature encoder is trained to extract features that are both semantically important and sparse. Second, a lightweight Feature Utility Estimator (FUE) head is introduced to predict the data utility of each agent’s feature to the final perception result, producing compact meta utility maps that (i) approximate the marginal contribution of each agent’s features in the BEV grid to the downstream task and (ii) expose cross-agent feature overlap. This entire framework is trained end-to-end, using a differentiable scheduler to jointly optimize the feature encoder and the FUE. During inference, agents exchange the meta utility maps and locally compute an explainable, deterministic transmission policy: select at most one agent with the highest utility per cell, and greedily admit cells until the communication budget is met, which is provably optimal under our optimization formulation.
@@
JigsawComm’s transmission policy inherently eliminates cross-agent duplication by construction with its top-$1$ per-cell rule. This makes the system practical and scalable: because the final transmission is a sparse ‘jigsaw’ assembled from the best cell-level features across the network, the total communication cost is bounded by the size of a single sparse feature map. This achieves an asymptotic transmission cost of $\mathcal{O}(1)$ with respect to the number of agents $N$, rather than the $\mathcal{O}(N)$ of naive broadcast. The scheduler requires only the exchanged meta utility maps and yields identical decisions across all agents, making it deployable even in the challenging decentralized mode of V2X.
Our contributions can be summarized as follows:
@@
  • We propose JigsawComm, a novel communication-efficient CP framework that maximizes the contribution of each transmitted bit to the final perception task, using a jointly trained semantic sparse feature encoder and end-to-end differentiable scheduler.
@@
  • We introduce a learned utility proxy and a deterministic, provably optimal top-1 scheduling policy that, effectively eliminates cross-agent redundancy, achieving scalable $\mathcal{O}(1)$ communication cost.
  • We extensively evaluate JigsawComm on the OPV2V and DAIR-V2X benchmarks, demonstrating comparable or superior accuracy while reducing communication data volume and improving utility-per-byte both by up to over $500\times$ compared to prior methods [xu2022opv2v, f_cooper, xu2022v2x, hu2022where2comm], setting the new SOTA for communication-efficient CP. [^1] Code available at https://github.com/WiSeR-Lab/JigsawComm.
@@ 2.2 Communication-Efficient CP Strategies
Message Representation: Changing How Information is Encoded. This category of methods focuses on compressing the feature maps. While standard encoder-decoder models are common [xu2022opv2v, xu2025cosdh], more advanced strategies have emerged. CodeFilling [hu2024communication, zhao2025quantv2x], for example, uses a shared codebook (i.e. lookup table) to quantize high-dimensional features into compact integer indices. A key limitation of this approach is that the codebook is learned offline and remains static during inference. This makes them difficult to generalize to novel environments. More recently, DiffCP proposed a generative approach [mao2025diffcp], where the sender transmits a small semantic seed, and the receiver leverages a diffusion model to reconstruct the sender’s full BEV map. While achieving very low bitrates, this method incurs a substantial computational burden for the iterative denoising process, violating real-time latency constraints [ADConstraints], and carries an inherent risk of “hallucination”, sacrificing accuracy.
@@
Message Selection: Deciding What to Transmit. This line of research aims to prune the information sent by agents, assuming that not all data is equally valuable. Early work like When2com [liu2020when2com] employed a graph neural network with a handshake mechanism to select which agents to communicate with, performing a coarse, agent-level selection. However, as V2X communication primarily relies on wireless broadcast rather than direct pairwise communications [FCC_DSRC, CV2X_ETSI_EN_303_613], optimizing communication graphs does not save bandwidth in practice. A more fine-grained approach was introduced by Where2comm [hu2022where2comm], which has become the de facto canonical baseline for communication-efficient CP. The idea is to generate a spatial confidence map by applying the detector head on each agent’s local feature. This map serves as a heuristic to identify ‘perceptually critical’ regions for transmission, which achieves region-level sparsity, reducing bandwidth compared to sending dense feature maps. More recent work followed the same idea (e.g[xu2025cosdh] further considered intermediate-late hybrid fusion). However, these methods still rely on heuristic utility proxies, which may not be optimal for the final CP task. They also lack explicit mechanisms to suppress redundant information from multiple collaborators with overlapping observations, resulting in a non-scalable, linear increase of communication cost with respect to the number of agents.
Nevertheless, quantifying and minimizing inter-agent redundancy is a challenging problem. Prior CP schemes that consider redundancy reduction suffer from various limitations. For example, a few early fusion based CP methods adopt geometry-based heuristics (e.g. each agent responsible for partitioned areas that are closest to itself) [zhang2021emp, autocast, wang2024edge, yu2023flow], which ignore sensory data content and semantics (e.g. occlusion), leading to potentially inaccurate results. Also, intermediate fusion-based CP methods that adopt mutual information-based metrics for redundancy reduction are computationally impractical with limited samples [su2024makes].
Table 1: Comparison with existing methods.
[…]
    Refers to whether the method can meet the bandwidth constraint of V2X and is scalable.
  • $\dagger$
    Refers to a method’s ability to adapt to variations in available bandwidth.
Positioning Our Work. The existing landscape of communication-efficient CP reveals a clear set of unresolved challenges: a reliance on heuristic and non-task-optimal proxies for task utility, a lack of explicit modeling and removal of semantic inter-agent redundancy, and static performance-bandwidth trade-offs imposed by fixed representation schemes. Our work addresses these gaps by introducing a semantic-aware, task-oriented, and end-to-end learned scheduling policy that is generalizable, adaptive, and efficiently manages shared communication resources among multiple agents. Table 1 provides a comparative summary.
Figure 2: Overview of JigsawComm.
3 Methodology: End-to-End Communication-Aware Cooperative Perception
@@ 3.1 Problem Formulation
We consider $N$ agents $\mathcal{A}=\{a_{1},\dots,a_{N}\}$ (CAVs or RSUs, within each other’s communication range) collaborating via V2X. Per cycle, agent $a_{i}$ obtains raw sensor data $X_{i}$ and uses an encoder $\Phi_{\text{enc}}$ to produce a BEV feature map $\mathbf{F}_{i}=\Phi_{\text{enc}}(X_{i})\in\mathbb{R}^{C\times H\times W}$ with cells $\mathbf{f}_{i}^{l}\in\mathbb{R}^{C}$, $l\in\mathcal{L}=\{1,\dots,L\}$, L=HWL{=}HW. After exchanging meta utility maps ${\mathbf{U}}_{i=1}^{N}$, each agent $a_{i}$ then transmits a sparse subset $\mathbf{F}^{\prime}_{i}=\mathbf{M}_{i}\odot\mathbf{F}_{i}$, where $\mathbf{M}_{i}\in\{0,1\}^{H\times W}$ and $\odot$ denotes element-wise multiplication. A fusion module $\Phi_{\text{fuse}}$ and decoder $\Phi_{\text{dec}}$ produce the output $\hat{Y}$. Our goal is to maximize the CP accuracy within a total communication budget BB [^2] In practice, $B$ equals the available bandwidth $BW$ divided by the frame rate $r$ (e.g. if $BW=20$ Mbps and $r=10$ FPS, $B=2$ Mb).:
@@
Here, $\mathcal{P}(\cdot)$ is the CP accuracy evaluation metric (e.g. mAP), $Y$ corresponds to the ground truth supervision, $\theta$ denotes the trainable parameters for the entire network, and the binary masks $\{\mathbf{M}_{i}\}_{i=1}^{N}$ select subsets of the encoded features across spatial dimensions for transmission and fusion.
Challenges. This formulation presents three core challenges: (C1) During inference, since there is no ground truth, the CP accuracy $\mathcal{P}$ cannot be directly evaluated by the agents. Thus, we must design a utility proxy that is aligned with the CP accuracy and can be computed locally. Previous work uses heuristic local confidence scores [liu2020when2com, hu2022where2comm, sheng2024semantic, xu2025cosdh], which fails to capture cross-agent redundancy and is not necessarily the optimal representations of the final accuracy. (C2) Ideally, a utility proxy needs to be jointly trained with the network under the accuracy objective to guarantee a direct alignment. However, the binary scheduling masks are non-differentiable, which hinders end-to-end training. (C3) Even if the utility function can be computed or approximated, the feature selection process under Eq. (2) is combinatorial, as it reduces to a 0-1 knapsack problem, which is NP-Hard [kellerer2004multidimensional]. Moreover, ‘more bits’ is not better: redundant views congest the V2X data channel without boosting fused utility. Thus, we must transmit only semantically meaningful, non-redundant information and maximize the contribution of every transmitted bit to CP accuracy.
System Overview. JigsawComm learns to assemble the ‘jigsaw’ of multi-agent BEV features so that the CP messages carry only essential, non-overlapping content. As shown in Fig. 2, each agent: 1 produces sparse and semantically important features (Sec. 3.2); 2 predicts a compact meta utility map using a Feature Utility Estimator (FUE) network, which is exchanged among agents efficiently (Sec. 3.3); 3 locally applies a deterministic, redundancy-aware top-1-per-cell policy, producing a feature selection mask that maximizes the total utility while satisfying the communication budget (Sec. 3.4); 4 broadcasts the selected features; 5 fuses the exchanged features and passes to the inference head for output. With the total per-frame communication cost upper-bounded by the size of a single sparse semantic feature, the required bandwidth is $\mathcal{O}(1)$ as the number of agents increases, ensuring high scalability.
@@ 3.2 Sparse Semantic Feature Encoder
To ensure that features carry compact and essential semantics using a minimal number of bits, we first shape the encoder’s output to reduce intra-agent redundancy. We apply a learnable threshold $\kappa$ and a semantic loss consisting of an $L_{1}$ regularization on the encoded feature:
@@
The $L_{1}$ regularization, when applied to model parameters, is known to produce sparse coefficients, suppress irrelevant features, and shape the coefficient amplitudes based on their quantitative contribution to the final outcome [ng2004feature]. However, applying $L_{1}$ directly to raw activations $\|\mathbf{f}_{i}^{l}\|_{1}$ only encourages small magnitudes, yet under stochastic normalization and noise, almost never yields exact zeros; gradients remain nonzero almost everywhere, so uninformative activations linger at small but nonzero values—still consuming bits. Therefore, we incorporate the learnable threshold $\kappa$, which explicitly introduces sparsity.

### v2

@@ 1 Introduction
Figure 1: A sample scenario from the OPV2V dataset (left). Under the same bandwidth limit, prior work [hu2022where2comm] (middle) does not utilize bandwidth efficiently, leading to missed detections. JigsawComm (right) maximizes every bit’s contribution, enabling complete and accurate detection. See Sec. 4 for quantitative evaluation and discussion on the “cost” of redundancy.
@@
Despite this promise, CP faces a formidable challenge in communication efficiency. The available V2X bandwidth is severely limited (up to $20$ MHz for both DSRC [FCC_DSRC] and C-V2X [CV2X_ETSI_EN_303_613]), while the volume of sensor data generated is prohibitively large (e.g. 4 MB per LiDAR frame [f_cooper]). Exchanging raw sensor data is therefore infeasible while satisfying the stringent real-time perception requirements of autonomous driving (e.g. <<100 ms [ADConstraints]). As previous research also demonstrated that exchanging local detection results alone is suboptimal [xu2022opv2v], the field has largely converged on the intermediate fusion framework—exchanging and fusing Bird’s-Eye-View (BEV) features—as the most effective paradigm [xu2022opv2v, hu2022where2comm, xu2022v2x, f_cooper]. Even so, transmitting dense deep features from multiple agents still exhausts V2X bandwidth.
@@
The key challenge stems from the high degree of data redundancy in the shared features, which takes two forms: (i) intra-agent redundancy, where features waste bits on semantically irrelevant information; and (ii) inter-agent redundancy, where multiple agents transmit overlapping data about the same scene regions. As illustrated in Fig. 1, under limited bandwidth such inefficiency not only reduces detection coverage but can even degrade perception accuracy when lower-quality data dilutes higher-quality representations after fusion (Sec. 4.4).
@@
Existing communication-efficient CP methods fail to address both forms simultaneously. They typically adopt one of two strategies: (i) Compression, using encoder-decoders [xu2022opv2v] or codebooks [hu2024communication, zhao2025quantv2x], which reduce data size but still transmit the entire feature map or restrict representation power to offline-trained dictionaries; (ii) Message selection [liu2020when2com, hu2022where2comm, hu2024communication, xu2025cosdh], choosing “critical” regions based on local heuristics (e.g. detection confidence) that loosely correlate with the final cooperative task. Crucially, both approaches ignore cross-agent redundancy, so their communication cost still grows linearly with the number of agents. The problem of designing a truly communication-efficient, scalable, and accurate CP system remains open.
@@
We argue that a practical CP system must maximize the contribution of every transmitted bit to the final perception goal, transmitting only essential, semantically meaningful, and non-redundant information. This requires jointly designing the feature encoding and transmission selection, which is non-trivial: removing cross-agent redundancy is a “chicken-and-egg” problem, since an agent cannot determine whether its features are redundant without knowing what other agents intend to transmit.
@@
To address these gaps, we propose JigsawComm, an end-to-end semantic-aware CP framework that learns to “assemble the puzzle” of multi-agent feature transmission. First, a regularized feature encoder extracts features that are both semantically important and sparse. Second, a lightweight Feature Utility Estimator (FUE) predicts each agent’s per-cell contribution to the final perception result, producing compact meta utility maps that approximate the marginal value of each feature and expose cross-agent overlap. The entire framework is trained end-to-end via a differentiable scheduler that jointly optimizes the encoder and FUE. At inference, agents exchange meta utility maps and locally compute an explainable, deterministic transmission policy: select at most one agent with the highest utility per cell, then greedily admit cells until the communication budget is met. We prove this policy is optimal under the learned utility proxy (Theorems 3.1 and 3.2) and validate its alignment with task accuracy empirically.
@@
The top-1-per-cell rule inherently eliminates cross-agent duplication by construction, making the system practical and scalable: the final transmission is a sparse “jigsaw” assembled from the best cell-level features across agents, bounding the data-channel cost to $\mathcal{O}(1)$ with respect to the number of agents $N$, compared to $\mathcal{O}(N)$ for naive broadcast. Although exchanging the lightweight meta utility maps incurs $\mathcal{O}(N)$ signaling, this overhead is negligible in practice (<<0.03 KB per frame). The scheduler yields identical decisions on every agent, making it deployable in both centralized (V2I) and decentralized (V2V) modes. Our contributions are:
@@
  • We propose JigsawComm, a CP framework that jointly trains a sparse semantic encoder and an end-to-end differentiable scheduler to maximize each transmitted bit’s contribution to perception accuracy.
@@
  • We introduce a learned, task-aligned utility proxy and a top-1-per-cell scheduling policy that eliminates cross-agent redundancy and bounds data-channel cost to $\mathcal{O}(1)$ with respect to the number of agents.
[…]
@@ 2.2 Communication-Efficient CP Strategies
Message Representation. This category compresses feature maps via encoder-decoders[xu2022opv2v, xu2025cosdh, Zhang_2024_CVPR], codebook quantization [hu2024communication, zhao2025quantv2x], or generative reconstruction [Zhang_2024_CVPR, mao2025diffcp, zhou2025pragmatic]. Encoder-decoder approaches apply a bottleneck to reduce the feature dimensionality, achieving fixed compression ratios but lacking adaptability to varying scene complexity. Codebooks approximate continuous features with discrete indices matching to an offline trained dictionary, which may restrict the generalizability of the representation [hu2024communication, zhao2025quantv2x]. Diffusion-based methods achieve high accuracy-bandwidth tradeoff (e.g. 0.86 mAP@0.5 under 80 Kbps [mao2025diffcp]) but the multi-step denoising violate real-time computation latency constraints for autonomous driving [ADConstraints]. Crucially, these approaches compress the entire feature map uniformly without considering which regions are task-relevant, wasting bits on uninformative background.
@@
Message Selection. This line of research aims to prune the information sent by agents, assuming that not all data is equally valuable. Early work like When2com [liu2020when2com] employed a GNN to select which agents to communicate with, performing a coarse, agent-level selection. However, as V2X communication primarily relies on wireless broadcast rather than direct pairwise communications [FCC_DSRC, CV2X_ETSI_EN_303_613], optimizing communication graphs does not save bandwidth in practice. A more fine-grained approach was introduced by Where2comm [hu2022where2comm], which has become the de facto canonical baseline for communication-efficient CP, which generate a spatial confidence map by applying the detector head on each agent’s local feature. This map serves as a heuristic to identify ‘perceptually critical’ regions for transmission, which achieves region-level sparsity, reducing bandwidth compared to sending dense feature maps. Follow-up work extends this idea with spatiotemporal fusion [tang2025cost] or intermediate-late hybrid fusion [xu2025cosdh]. However, these methods rely on local detection confidence that may not correlate well with the final CP accuracy and lack mechanisms to suppress inter-agent redundancy, resulting in $\mathcal{O}(N)$ communication scaling. Prior attempts at redundancy reduction either use geometry-based heuristics that ignore semantics [zhang2021emp, autocast, wang2024edge, yu2023flow] or mutual information metrics that are computationally impractical [su2024makes, wei2025infocom].
Positioning Our Work. Existing communication-efficient CP methods share three unresolved shortcomings: reliance on heuristic utility proxies; no explicit modeling of semantic inter-agent redundancy; and static performance-bandwidth trade-offs. Our work addresses all three by introducing a semantic-aware, task-oriented, end-to-end learned scheduling policy that is adaptive to varying bandwidth. While JigsawComm shares the high-level goal of task-oriented compression with semantic/goal-oriented communications [strinati2021_6g, pezone2022goal], existing work in that field targets single-source scenarios [yuan2023scalable, binucci2022adaptive] or neglects inter-sensor correlations [sheng2024semantic]. CP is inherently a multi-source problem where inter-agent redundancy must be modeled without knowing the data correlation structure a priori; JigsawComm solves this in a data-driven manner, with minimal overhead.
Figure 2: Overview of JigsawComm. Each agent encodes sparse BEV features, estimates a meta utility map via the FUE, exchanges utility maps, and applies the deterministic top-1-per-cell scheduler to select non-redundant features for transmission and fusion.
3 Methodology
@@ 3.1 Problem Formulation
We consider $N$ agents $\mathcal{A}=\{a_{1},\dots,a_{N}\}$ collaborating via V2X [^2] Here, agents can be CAVs or RSUs, within the same V2X communication range.. Per cycle, agent $a_{i}$ obtains raw sensor data $X_{i}$ and uses an encoder $\Phi_{\text{enc}}$ to produce a BEV feature map $\mathbf{F}_{i}=\Phi_{\text{enc}}(X_{i})\in\mathbb{R}^{C\times H\times W}$ with cells $\mathbf{f}_{i}^{l}\in\mathbb{R}^{C}$, $l\in\mathcal{L}=\{1,\dots,L\}$, L=HWL{=}HW. After exchanging meta utility maps ${\mathbf{U}}_{i=1}^{N}$, each agent $a_{i}$ then transmits a sparse subset $\mathbf{F}^{\prime}_{i}=\mathbf{M}_{i}\odot\mathbf{F}_{i}$, where $\mathbf{M}_{i}\in\{0,1\}^{H\times W}$ and $\odot$ denotes element-wise multiplication. A fusion module $\Phi_{\text{fuse}}$ and decoder $\Phi_{\text{dec}}$ produce the output $\hat{Y}$. Our goal is to maximize the CP accuracy within a total communication budget BB [^3] BB equals the available bandwidth $BW$ divided by the frame rate $r$.:
@@
Here, $\mathcal{P}(\cdot)$ is the CP accuracy evaluation metric (e.g. mAP), $Y$ corresponds to the ground truth supervision, $\theta$ denotes the trainable parameters for the entire network, and the binary masks $\{\mathbf{M}_{i}\}_{i=1}^{N}$ select subsets of the encoded features across spatial dimensions for transmission and fusion. During training, Eq. (1) is optimized jointly over the model parameters $\theta$ and the transmission masks $\{\mathbf{M}_{i}\}$ via the differentiable scheduler (Sec. 3.4); at inference, $\theta$ is fixed and only the scheduling masks $\{\mathbf{M}_{i}\}$ are optimized.
Challenges. This formulation presents three core challenges: (C1) During inference, each agent must decide which features to transmit without access to ground truth or full features from other agents (infeasible to transmit) to optimize for $\mathcal{P}(\cdot)$. Thus, a locally computable and compact utility proxy that correlates with CP accuracy and captures cross-agent redundancy is required. (C2) The binary scheduling masks $\mathbf{M}_{i}$ are non-differentiable, preventing direct end-to-end training of the utility proxy. (C3) The feature selection under Eq. (2) is a 0-1 knapsack problem, which is NP-Hard in general [kellerer2004multidimensional].
System Overview. JigsawComm learns to assemble the ‘jigsaw’ of multi-agent BEV features so that the CP messages carry only essential, non-overlapping content. As shown in Fig. 2, each agent: 1 produces sparse and semantically important features (Sec. 3.2); 2 predicts a compact meta utility map using a Feature Utility Estimator (FUE) network, which is exchanged among agents efficiently (Sec. 3.3); 3 locally applies a deterministic, redundancy-aware top-1-per-cell policy, producing a feature selection mask that maximizes the total utility while satisfying the communication budget (Sec. 3.4); 4 broadcasts the selected features; 5 fuses the exchanged features and passes to the inference head for output. With the total per-frame data payload upper-bounded by the size of a single sparse semantic feature, the data-channel bandwidth scales as $\mathcal{O}(1)$ with respect to the number of agents, ensuring high scalability.
@@ 3.2 Sparse Semantic Feature Encoder
To ensure that features carry compact and essential semantics using a minimal number of bits, we first shape the encoder’s output to reduce intra-agent redundancy. We apply a semantic loss consisting of an $L_{1}$ regularization and a learnable threshold $\kappa$ on the encoded feature $\mathbf{f}_{i}$:
@@
$L_{1}$ regularization is known to produce sparse coefficients and suppress irrelevant features [ng2004feature]. However, applying it directly to raw activations ($\|\mathbf{f}_{i}^{l}\|_{1}$) only encourages small magnitudes without yielding exact zeros—uninformative activations linger at small but nonzero values, still consuming bits. The learnable threshold $\kappa$ resolves this by explicitly zeroing out activations below $\kappa$, producing structurally sparse features with exact zeros that incur no transmission cost under sparse index encoding [hu2022where2comm].